You buy a 5-year zero coupon bond with 4% yield to maturity. You sell the bond 2 years later when it's yield to maturity is 2%. What was your annualized holding period return?
Answer #1

Assuming a par value of $1000,

Price of the bond at 5 year is- =1000*1.04^-5 = 821.927

Price of the bond 3 years before maturity  =1000*1.04^-3 =888.996

Holding period return =888.996/821.927-1 = 8.16%